ULTRAFAST EXCITED-STATE DYNAMICS IN MODEL HEXABROMOPLATINATE (IV) AND HEXABROMOOSMIATE (IV) DIANIONS IN THE CONDENSED PHASE

\underline{I. L. ZHELDAKOV}, A. S. MERESHCHENKO, A. N. TARNOVSKY, Bowling Green University, Department of Chemistry and Center for Photochemical Sciences, Bowling Green, Ohio 43403.

Photoexcitation of hexabromoplatinate and hexabromoosmate dianions in CT and ligand field states results in remarkably fast formation (sub 100 fs) of coherently excited photoproducts assigned to ^{3}P tBr_{5}^{−} and ^{3}OsBr_{5}^{−}. This is explained by the fact that one of the lowest triplet excited states in both P tBr_{6}^{2−} and OsBr_{6}^{2−} is repulsive. Surprisingly, the damping time and amplitude of the observed coherent oscillations in the pentabromoosmate (209 cm^{−1}) and pentabromoplatinate (112 cm^{−1}) species is quite different. The shorter damping time (0.4 ps) of ^{3}P tBr_{5}^{−} in comparison with the damping time (1 ps) of ^{3}OsBr_{5}^{−} in aqueous solution is related to faster dephasing in ^{3}P tBr_{5}^{−}. The large-amplitude bending oscillations in pentabromoplatinate are tentatively assigned to the passage of the wavepacket through the Janh-Teller conical intersection(s), whereas in pentabromoosmate, the umbrella type of oscillations appears to be due to the coupling to the Os-Br symmetric stretching mode in the parent dianion. The coherent oscillations in both photoproducts show no dependence on solvent. The results are supported by DFT, TD-DFT and CASSCF/CASPT2 calculations of electronic structures, vertical electronic transitions and harmonic frequencies of initial dianions and photoproducts.